a man's age is three times his son's age. Ten years ago he was five times his son's age. Calculate the age of each person
Answer by JBarnum(2146)

M=3S
M-10=5(S-10)
M-10=5S-50
3S-10=5S-50
40=2S
20=S
M=3S
M=3(20)
M=60
Man 60
Son 20